# Quantum Convolutional Neural Networks for Jet Images Classification

### v1

@@ Abstract
Recently, interest in quantum computing has significantly increased, driven by its potential advantages over classical techniques. Quantum machine learning (QML) exemplifies one of the important quantum computing applications that are expected to surpass classical machine learning in a wide range of instances. This paper addresses the performance of QML in the context of high-energy physics (HEP). As an example, we focus on the top-quark tagging, for which classical convolutional neural networks (CNNs) have been effective but fall short in accuracy when dealing with highly energetic jet images. In this paper, we use a quantum convolutional neural network (QCNN) for this task and compare its performance with CNN using a classical noiseless simulator. We compare various setups for the QCNN, varying the convolutional circuit, type of encoding, loss function, and batch sizes. For every quantum setup, we design a similar setup to the corresponding classical model for a fair comparison. Our results indicate that QCNN with proper setups tend to perform better than their CNN counterparts, particularly when the convolution block has a lower number of parameters. This suggests that quantum models, especially with appropriate encodings, can hold potential promise for enhancing performance in HEP tasks such as top quark jet tagging.
@@ I Introduction
Classical convolutional neural networks (CNNs) have been widely and successfully employed for top-quark tagging [15]. However, they struggle to provide the required accuracy when faced with the highly energetic and complex top-quark jet images [16]. Aiming to overcome this problem, our main objective is to identify a practically suitable quantum machine learning architecture to classify top-quark and QCD jet images. We do this by quantifying its accuracy and comparing its performance with a classical machine learning method. Specifically, we focus on a quantum convolutional neural network (QCNN), first introduced by Cong et al. [17], which uses shallow-depth quantum circuits, making them suitable for the noisy intermediate-scale quantum (NISQ) devices that are already currently accessible. This shallowness also yields robustness against barren plateau issues which guarantees the trainability of its circuit [18]. Also, results from previous works that apply the QCNN to classical images indicated improved accuracy for QCNN over CNN [19, 20] under certain conditions, possibly due to the entangling property in a quantum model.
@@
In this study, we utilize publicly available JetNet library datasets [21], specifically the TopTagging dataset which contains top and QCD jets. The top jets in this dataset are in the hadronic channel. The current stage of this study is to have a proof of principle that our quantum neural network can be utilized in HEP analysis such as in top-tagging. Hence, this work’s analysis will involve the top decay’s hadronic channel rather than the leptonic channel, since it’s the provided channel in the TopTagging dataset. However, from the jet images pre-processing we could already notice the similarity of top and QCD images which yields the required complexity to test our quantum and classical models. After forming the jet images, the principal component analysis (PCA) is applied to reduce the dimensionality of the formed images. The quantum model is implemented on a classical noiseless simulator (PennyLane [22]), while the classical counterpart is built using the TensorFlow framework [23]. Since we are working with a classical dataset, incorporating an appropriate encoding in our quantum circuit model is essential. So we compared various setups for the QCNN, varying the convolutional circuit, type of encoding, loss function, and batch sizes. We design a similar setup for every quantum setup to the corresponding classical model and use the resulting performance as a reference.
@@
The structure of this paper goes as follows. Section II describes the methodology of our study. In this section, we start by introducing the main components of a basic classical CNN model, then we explain how these concepts can be implemented quantum mechanically to build a QCNN. After that, we discuss the formation and the pre-processing of the jet images. The results and QCNN to CNN comparisons are presented in section III. Then, we conclude and mention future directions and further plans of our study in section IV.
@@ II.3.2 Convolutional and Pooling circuits
After that comes the pooling layer to reduce the dimensionality of the system by half. The pooling circuit, shown in Fig. 4c, comprises a CNOT gate and 9 single-qubit gates all containing trainable parameters. The convolutional and pooling layers are repeated until the system has reached a sufficiently small size, one qubit in our case in which its output is measured with a PauliZ measurement, and the predicted output is then compared to the true label for the classification task as explained in Section II.2.
@@
II.4 Dataset Preparation
@@
where its components are referred to as the particle’s features. The maximum number of particles provided in each jet is 200. The dataset is then preprocessed and transformed to form images. Then, we apply PCA to reduce dimensionality of the images. This reduction enables the use of a simple quantum circuit where each qubit takes a single pixel value from the image as an input.
II.4.1 Jet images
@@ III.1 Dependence on Encodings
Table 1: Test accuracy for tensor product embedding (TPE), hardware efficient embedding with one later (HEE1) and two layers (HEE2), and the classically hard embedding (CHE) results averaged over 50 runs for $SO(4)$ with 30 parameters and CNN results with 33 parameters, and $SU(4)$ with 48 parameters and the comparable CNN with 51 parameters. The batch-size is fixed to 32.
@@
Accuracy results after $30$ epochs averaged over $50$ runs for QCNN with $SO(4)$ and $SU(4)$ with varying (quantum) encoding types compared to CNN for the three loss functions are presented in Table 1. Here, the batch-size is fixed to 32. We can notice that accuracy results for QCNN are better than those of CNN when we use $SO(4)$ along with Hinge (except for CHE encoding), MSE, or cross-entropy. The performance of QCNN for $SU(4)$ setups is similar or worse. We can also observe the relatively large error for CNN results at the low parameter regime. For QCNN setups, first with relatively high accuracy, HEE1 encoding (values given in bold in Table 1), and then TPE showed better results compared to HEE2 or CHE. This could be possibly understood from dataset induced barren plateau investigated in [25]. In other words, the HEE2 and CHE encoding would be too complex specifically for this task, which induces trainability issues (barren plateau). However, we have to study performance by increasing the number of qubits to make this connection more concretely, which we leave for the future direction.
@@ IV Conclusion and outlook
In this paper, we utilized two QCNN architectures one with $SO(4)$ convolutional circuit and the other built with an $SU(4)$ convolutional circuit. We varied the setups of each architecture to test different loss functions, encodings, and batch-sizes. We compared each setup with a corresponding CNN model. From the preliminary results we got for our quantum and classical models, we found that the QCNN, in the absence of noise, performs better than the corresponding CNN model for a specific choice of data encoding, batch-size, loss function, and convolutional block, at least for this task along with PCA, but do not otherwise. However, we need to emphasize that the system we studied here was relatively simple, both in QCNN and CNN, therefore, it was possible to train the classical model with a very small number of parameters, whereas the usual case of training CNN models is to use a high number of parameters that may reach up to millions. Besides, we only performed the noiseless simulation and the effects from noise would spoil the QCNN performance unless we can tame them.
Aiming to noisy hardware implementation, it would be better to reduce the complexity of a circuit as much as possible. We could try the following two methods for this purpose. The first one is dimensional expressivity analysis (DEA) proposed in [30]. DEA helps detect redundant parameters and potentially optimize the circuit which may result in reduced training time, less noise from quantum gates, and better suitability for real quantum hardware. Current DEA is done after setting up all circuit elements including embedding gates. Because embedding circuits in QML can change depending on different input images, we have to modify DEA to check for the redundant parameters for every single input image. If this is achieved successfully, it might result in a dynamic QNN (DQNN) which could be more flexible and adaptive to the type of data used. The second one is the equivariant QNN (EQNN) [31] which mainly relies on the symmetry found in data. EQNN methods allow us to construct QNNs that automatically satisfy the desired symmetry constraints. Since our dataset can also have rotational symmetry thus potentially, we can apply these techniques.
On top of reducing the circuit complexity, an important ingredient for the realization on noisy devices is an error mitigation method, such as zero noise extrapolation (ZNE) [32, 33, 34], probabilistic error cancellation [33, 35], and Twirled Readout Error eXtinction (TREX) [36].
Ultimately, it would be important to study the scaling of performance and resources with respect to the complexity of the dataset and/or the number of qubits. It would be interesting if we could combine the techniques listed above to see the scaling on the actual hardware.

### v2

@@ Abstract
Recently, interest in quantum computing has significantly increased, driven by its potential advantages over classical techniques. Quantum machine learning (QML) exemplifies one of the important quantum computing applications that are expected to surpass classical machine learning in a wide range of instances. This paper addresses the performance of QML in the context of high-energy physics (HEP). As an example, we focus on the top-quark tagging, for which classical convolutional neural networks (CNNs) have been effective but fall short in accuracy when dealing with highly energetic jet images. In this paper, we use a quantum convolutional neural network (QCNN) for this task and compare its performance with CNN using a classical noiseless simulator. We compare various setups for the QCNN, varying the convolutional circuit, type of encoding, loss function, and batch sizes. For every quantum setup, we design a similar setup to the corresponding classical model for a fair comparison. Our results indicate that QCNN with proper setups tend to perform better than their CNN counterparts, particularly when the convolution block has a lower number of parameters. For the higher parameter regime, the QCNN circuit was adjusted according to the dimensional expressivity analysis (DEA) to lower the parameter count while preserving its optimal structure. The DEA circuit demonstrated improved results over the comparable classical CNN model.
@@ I Introduction
Classical convolutional neural networks (CNNs) have been widely and successfully employed for top-quark tagging [15]. However, they struggle to provide the required accuracy when faced with the highly energetic and complex top-quark jet images [16]. Aiming to overcome this problem, our main objective is to identify a practically suitable quantum machine learning architecture to classify top-quark and QCD jet images. We do this by quantifying its accuracy and comparing its performance with a classical machine learning method. Specifically, we focus on the quantum convolutional neural network (QCNN), first introduced by Cong et al. [17], which leverages shallow-depth quantum circuits, making it suitable for the noisy intermediate-scale quantum (NISQ) devices currently accessible. This circuit shallowness also provides robustness against barren plateau issues, ensuring the trainability of the model [18]. While these advantages might make QCNN classically simulable, as highlighted in [19], the classical simulation of QCNN can still be regarded as a quantum-inspired machine learning architecture. In fact, several previous works have demonstrated improved accuracy of QCNN over conventional classical CNNs under certain conditions [20, 21].
@@
In this study, we utilize publicly available JetNet library datasets [22], specifically the TopTagging dataset which contains top and QCD jets. The top jets in this dataset are in the hadronic channel. The current stage of this study is to have a proof of principle that our quantum neural network can be utilized in HEP analysis such as in top-tagging. Hence, this work’s analysis will involve the top decay’s hadronic channel rather than the leptonic channel, since it’s the provided channel in the TopTagging dataset. However, from the jet images pre-processing we could already notice the similarity of top and QCD images which yields the required complexity to test our quantum and classical models. After forming the jet images, the principal component analysis (PCA) is applied to reduce the dimensionality of the formed images. The quantum model is implemented on a classical noiseless simulator (PennyLane [23]), while the classical counterpart is built using the TensorFlow framework [24]. Since we are working with a classical dataset, incorporating an appropriate encoding in our quantum circuit model is essential. So we compare various setups for the QCNN, varying the convolutional circuit, type of encoding, loss function, and batch sizes. To achieve an optimal QCNN structure, we conduct a dimensional expressivity analysis (DEA) to remove less significant gates from the circuit while preserving maximal expressivity. We design a similar setup for every quantum setup to the corresponding classical model and use the resulting performance as a reference.
@@
The structure of this paper goes as follows. Section II describes the methodology of our study. In this section, we start by introducing the main components of a basic classical CNN model. We then explain how these concepts can be translated into the quantum domain to construct a QCNN, which is subsequently analyzed using DEA. After that, we discuss the formation and the pre-processing of the jet images. The results and QCNN to CNN comparisons are presented in section III. Then, we conclude and mention future directions and further plans of our study in section IV.
@@ II.2 Loss functions
Finally, these loss functions are optimized via the Adam optimizer [26] in this work. This method combines several algorithms such as momentum and RMSProp optimizers to obtain reliable learning results.
@@ II.3.2 Convolutional and Pooling circuits
After that comes the pooling layer to reduce the dimensionality of the system by half. The pooling circuit, shown in Fig. 4c, comprises a CNOT gate and 9 single-qubit gates all containing trainable parameters. The convolutional and pooling layers are repeated until the system has reached a sufficiently small size, one qubit in our case. The output of this qubit is measured with a PauliZ measurement. The prediction of the neural network corresponds to the expectation value of the PauliZ operator and is then compared to the true label for the classification task as explained in Section II.2.
@@
II.4 DEA
Dimensional expressivity analysis (DEA), introduced by Funcke et al. in [30] and [31], is a technique designed to optimize a parametric quantum circuit (PQC) by achieving maximal expressivity with the fewest trainable parameters. In other words, it enables the construction of a PQC that retains its full expressive capability while minimizing redundancy in its parameters. Maximal expressivity here refers to the ability of the circuit to generate all relevant physical states, while a minimal set of parameters ensures that only independent trainable parameters remain, with any redundant ones either eliminated or set constant.
To formalize this, let us consider the PQC as a map $C$ that takes the parameters from the parameter space $\mathcal{P}$ to the state space $\mathcal{S}$. For any $\theta\ \epsilon\ \mathcal{P}$, the output $C(\theta)$ would be a state generated by the quantum circuit.
To identify redundant parameters, we first check the derivative of the PQC map with respect to each parameter $\theta_{k}$, denoted as $\partial_{k}C(\theta)$. If $\theta_{k}$ is redundant, then $\partial_{k}C(\theta)$ can be expressed as a linear combination of the remaining partial derivatives $\partial_{j}C(\theta)$ for $j\neq k$. A clear way to see this is through the Jacobian $J$ of $C$
$J_{k}(\theta)=\begin{pmatrix}\mathcal{R}\partial_{1}C(\theta)&...&\mathcal{R}\partial_{k}C(\theta)\\ \mathcal{I}\partial_{1}C(\theta)&...&\mathcal{I}\partial_{k}C(\theta)\end{pmatrix},$ (5)
where we check the rank of the $J$ inductively for each $k$. If adding a column $\partial_{k}C(\theta)$ does not increase the rank of $J$, then $\theta_{k}$ is deemed redundant and can be eliminated. Once the rank of $J$ found equal the dimension of $\mathcal{S}$, all remaining parameters are redundant and no further checks are necessary.
Another efficient way to identify redundant parameters is through the $S$-matrix, which is defined as
$S_{k}=J_{k}^{*}J_{k}.$ (6)
The matrix $J_{k}$ has a full rank, if and only if all eigenvalues of $S_{k}$ are positive. If, however, the smallest eigenvalue of $S_{k}$ was found to be zero, then we know that $\theta_{k}$ is redundant.
II.5 Dataset Preparation
@@
where its components are referred to as the particle’s features. The maximum number of particles provided in each jet is 200. The dataset is then preprocessed and transformed to form images. Then, we apply PCA to reduce the dimensionality of the images. This reduction enables the use of a simple quantum circuit where each qubit takes a single pixel value from the image as an input.
II.5.1 Jet images
@@ III.1 Dependence on Encodings
Accuracy results after $30$ epochs averaged over $50$ runs for QCNN with $SO(4)$ and $SU(4)$ with varying (quantum) encoding types compared to CNN for the three loss functions are presented in Table 1. Here, the batch-size is fixed to 32. We can notice that accuracy results for QCNN are better than those of CNN when we use $SO(4)$ along with Hinge (except for CHE encoding), MSE, or cross-entropy. The performance of QCNN for $SU(4)$ setups is similar or worse.
@@
Table 1: Test accuracy for Hinge (H), MSE (M), and Cross-entropy (C) loss function with the different encodings results averaged over 50 runs for $SO(4)$ with 30 parameters and CNN results with 33 parameters, and $SU(4)$ with 48 parameters and the comparable CNN with 51 parameters. The batch-size is fixed to 32.
We can also observe the relatively large error for CNN results at the low parameter regime. For QCNN setups, first with relatively high accuracy, HEE1 encoding (values given in bold in Table 1), and then TPE showed better results compared to HEE2 or CHE. This could be possibly understood from dataset induced barren plateau investigated in [27]. In other words, the HEE2 and CHE encoding would be too complex specifically for this task, which induces trainability issues (barren plateau). However, we have to study performance by increasing the number of qubits to make this connection more concrete, which we leave for the future direction.
@@ III.2 Dependence on Batch-sizes
III.3 DEA circuit
QCNN with the $SU(4)$ convolutional circuit demonstrated better performance compared to the $SO(4)$ convolutional circuit. However, it required more parameters and, hence, longer training times. Now the question is, will we get similar results if we train the QCNN $SU(4)$ circuit after removing the redundant parameters using DEA?
For this task, we select the setup that resulted in the best accuracies for both QCNNs and CNNs. This setup, as shown in Table 1, corresponds to the use of MSE loss along with the HEE1 encoding. After performing a DEA check on the $SU(4)$ QCNN circuit with originally 48 parameters, 17 of them were found to be redundant, leaving a total of 31 parameters in the optimized circuit. This is compatible with the dimension of the state space ($12^{Q+1}-$ 1, where $Q$ is the number of qubits in the system). Therefore, the circuit is minimal in the set of trainable parameters and maximally expressive.
Test accuracy results of the MSE loss averaged over 1000 runs for $SU(4)$ and $SO(4)$ QCNN circuits before and after applying DEA, compared to CNN accuracy can be found in Table 3 and Fig. 8. The accuracy values at epoch 30 in Table 3 indicate a clear difference between the case when we have comparable number of parameters (31) of the DEA circuit with the CNN model (33 parameters) where the DEA circuit has higher accuracy value. The curves in Fig 8 also depict this result between all QCNN circuits (with and without DEA) compared to the CNN accuracy. We may notice slight differences between the $SU(4)$ circuit with and without DEA and the $SO(4)$ circuit at early epochs. The $SU(4)$ circuit without DEA has more parameters (48) which help it converge to the solution after fewer training epochs compared to the case with less parameters in the circuit with DEA (31). Both $SO(4)$ circuits, with and without DEA have a similar performance, and again we notice the slightly higher accuracy at early epochs for the circuit with the higher number of parameters. However, the $SU(4)$ circuit with DEA, has much less parameters compared to its original circuit and yet it shows comparable accuracy values. The $SO(4)$ circuit has 30 parameters, which is very close to the 31 parameters in the $SU(4)$ circuit with DEA. Notably, the DEA circuit shows comparable performances and even improved accuracy in the early epochs, suggesting that DEA can reduce the number of trainable parameters while maintaining strong performance.
[…]
$SU(4)$ without DEA 48 99.93 ± 0.01
$SU(4)$ with DEA 31 99.86 $\pm$ 0.01
$SO(4)$ without DEA 30 99.87 ± 0.01
$SO(4)$ with DEA 25 99.85 ± 0.01
CNN 33 93.08 ± 0.55
Table 3: Test accuracy results at epoch 30 averaged over 1000 runs using MSE loss function and HEE1 embedding with the various circuits studied.
Figure 8: Test accuracy results of the different circuits including and excluding DEA compared to CNN.
@@ IV Conclusion and outlook
In this paper, we utilized two QCNN architectures one with $SO(4)$ convolutional circuit and the other built with an $SU(4)$ convolutional circuit. We varied the setups of each architecture to test different loss functions, encodings, and batch-sizes. We compared each setup with a corresponding CNN model. From the preliminary results we got for our quantum and classical models, we found that the QCNN, in the absence of noise, performs better than the corresponding CNN model for specific choices of data encoding, batch-size, loss function, and convolutional block. However, it is important to emphasize that the system we studied here was relatively simple, both in QCNN and CNN. As a result, it was possible to train the classical model with a very small number of parameters, whereas the usual case of training CNN models is to use a high number of parameters that may reach up to millions. Besides, we only performed the noiseless simulation and the effects from noise would spoil the QCNN performance unless we can tame them.
In our study, DEA demonstrated the ability to significantly reduce the number of trainable parameters while maintaining good overall accuracy, thus enabling the QCNN to achieve higher accuracy compared to the CNN model with a similar parameter count. This could have a great impact when one is aiming for noisy hardware implementation, in which it will reduce the complexity of a circuit, and hence the computational complexity on the real device. Another way of reducing the complexity of the problem is through equivariant QNN (EQNN) [34] which mainly relies on the symmetry found in data. EQNN methods allow us to construct QNNs that automatically satisfy the desired symmetry constraints. Since our dataset can also have rotational symmetry thus potentially, we can apply these techniques.
As QCNN has been indicated to be classically simulable [19] for the classification of classical data, and at the same time it showed improved classification accuracy over the conventional classical CNN models for a number of case studies, one could use this as an advantage to study the scaling of performance and resources with respect to the complexity of the dataset and/or the number of qubits. The classical simulability helps in avoiding using real quantum hardware, which makes it possible to use a much higher number of qubits via classical simulation, and hence, for our problem of jet images classification, we can use higher dimensional images without going through PCA which is a more natural practice of dealing directly with pixels of the original dataset. Another interesting direction would be to apply this to quantum dataset which could also help in analysing experimental HEP problems such as studying the quantum properties of the final state particles detected from a collision of interest e.g. spin entanglement [35].